OpenStudy (anonymous):

Write the standard form of the line that passes through the given points. (7, -3) and (4, -8)

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@Nnesha

OpenStudy (phi):

I would first try to find the slope and y-intercept form, and then change to "standard form" to find the slope, find change in y divided by change in x

OpenStudy (anonymous):

Idk how to do that.

OpenStudy (phi):

you have two (x,y) pairs (the second number in the pair is the y value) change in y means find the difference in the y numbes can you do that ?

OpenStudy (anonymous):

..?

OpenStudy (phi):

in (7, -3) and (4, -8) can you list the 2 y values ?

OpenStudy (anonymous):

-3 and -8

OpenStudy (phi):

now find the difference : -3 - -8 = ?

OpenStudy (anonymous):

5

OpenStudy (phi):

that is the change in y now list the x values (in the same order as the y's) 7 and 4 fidn the difference : 7 - 4 = ?

OpenStudy (anonymous):

3

OpenStudy (phi):

now you know the slope: 5/3 now use the point slope form y - y1 = 5/3(x-x1) where (x1,y1) is a point on the line

OpenStudy (phi):

if we use the first point (7,-3) (we can use either point) y - -3 = 5/3(x - 7) or y + 3 = 5/3 x -35/3 standard form does not have fractions , so multiply both sides by 3 (that means both sides and all terms by 3)

OpenStudy (anonymous):

???